Calibration status for Run 14 production

Speaker : Gene Van Buren ( BNL )


Talk time : 12:35, Duration : 00:20
Run 14 AuAu200 calibrations status

Legend:
Not done
Done but not in database
Done and in database
  • EMC:
    • BTOW pedestal & status tables: done
  • TPC:
    • Alignment: done (presented at a recent S&C meeting)
    • T0: done as part of alignment (feedback from fine tuning confirmed)
    • Drift velocities: done (presented at recent S&C meeting, no feedback yet from fine tuning)
    • SpaceCharge & GridLeak: done (documentation)
      • No sector-by-sector GridLeak, no event-by-event corrections (fine tuning possibilities)
    • Hit errors: done
    • dE/dx: done
      • preliminary dE/dx was in database for the past ~week, but was not in the ongoing preview production
      • final dE/dx in database yesterday
  • HFT (IST & PXL):
    • Hit errors: done
    • Alignment: done (feedback from fine tuning resulted in ~100 μm shifts)
    • Status/masking tables: PXL is done, IST is done done (documentation for IST and PXL)
  • TOF:
    • Alignment: done (QA of local x and local y)
    • VPD slewing: done (single detector resolution of 68 ps, QA)
    • BTOF slewing: done
    • BTOF: zLocal: done
    • BTOF: T0: done
    • None of the above are in the database yet. Final checks ongoing and explected to complete ~today
    • Current findings are ~75-85 ps resolution for VPD+BTOF